# Tuning-Free Adaptive Style Incorporation for Structure-Consistent Text-Driven Style Transfer

### v1

@@ 5.1.2 Visual Enhancement for Real-World Images
Benefiting from the text-guided style transfer manner instead of the image-guided one, the diffusion-based method can be broadened to application areas unreachable by traditional image-to-image style transfer methods [58, 24, 14], particularly those related to visual enhancements such as vividness, realism, and depth of field. This is because these "styles" cannot be represented through an image.
As shown in Fig. 5, while P2P [19], PNP [43], and SDEdit [32] are based on the diffusion model, they only preserve the overall structure of the source image, sacrificing the detailed structure information for achieving effective visual enhance. We only present SDEdit with $0.6$ denoising strength, because excessive denoising strength usually causes drastic drifts in the image structure. This limits the application of the above methods to this specific task, as it necessitates strict consistency between the source and target images. Thanks to our proposed fine-grained style incorporation idea, our method efficiently enhances the visual quality of images while maintaining structural consistency. In terms of appearance enhancement, our method better maintains the horse’s body (Row. 1), the cat’s eyes (Row. 2), and the goldfish’s body details (Row. 3). In addition, our method achieves blurring under scene-invariant and semantic-invariant conditions, such as Spiderman’s left arm (Row. 4), and background details in Row. 5 and 6. These impressive results reveal the immense potential of diffusion models in addressing low-level vision tasks.
5.1.3 Style Transfer for Generated Images
As shown in Fig. 6, our method effectively transfers the source image to the target style while maintaining a high degree of structural consistency with the source image. Compared to our method, other methods are more prone to exhibiting structure deviations. As shown in Row. 1-3, the results of other methods can preserve the overall structure of the source image, yet they introduce notable modifications to the key visual features of the object. Furthermore, in Row. 4-6, the stylized images fail to maintain overall structural consistency with the source image. Our results achieve effective style transfer while minimizing modifications to the source image.
@@ 5.1.3 Style Transfer for Generated Images
5.2 Ablation Study
We conduct ablation experiments on the guidance mask of the AdaBlending module. As shown in Fig. 7, when both spatial level and attention-head level masks are simultaneously utilized (Col. 2), AdaBlending achieves consistent structural and semantic style transfer results. However, when only one guidance mask is used (Col. 3 and 4), noticeable structural drifts are observed in the girl’s head and eyebrow details. Furthermore, without using any guidance mask (Col. 5), additional semantic drifts are observed in the T-shirt color. These results highlight the importance of guidance masks in ensuring accurate and consistent style transfer outcomes.
6 Conclusion
In this work, we aim to study text-driven style transfer. To mitigate the well-known issue of structure inconsistency induced by text-image misalignment in diffusion-based T2I models, we propose a novel and effective tuning-free algorithm, termed Adaptive Style Incorporation (ASI), to incorporate the style information on the fine-grained feature level. It consists of the Siamese Cross-Attention (SiCA) to decouple the single-track cross-attention to a dual-track structure to obtain separate content and style features and the Adaptive Content-Style Blending (AdaBlending) module to couple the content and style information from a structure-consistent manner. We conduct a thorough evaluation of our method across diverse source image domains, both real and generated ones, on a range of common style transfer tasks. Experimental results have demonstrated the effectiveness of our proposed ASI and exhibit significant advantages over SOTA methods. Limitations and future work are provided in the Appendix.

### v2

@@ V-B2 Visual Enhancement
Benefiting from the text-guided style transfer method instead of the image-guided one, the diffusion-based method can be expanded to application areas unreachable by traditional image-to-image style transfer methods [65, 26, 15], particularly those related to visual enhancements such as realism, depth of field, and other photographic techniques. This is because these ”styles” cannot be represented through an image.
As shown in Fig. 7, while FreeDiff [55], P2P [20], PNP [50], and SDEdit [33] are based on the diffusion model, they only preserve the overall structure of the source image and cannot achieve effective visual enhancement according to the given prompt. This limits the application of the above methods to this specific task as it requires a more strict consistency between the source and target images.
Thanks to our proposed fine-grained style incorporation idea, our method efficiently enhances the visual quality of images while maintaining structural consistency. The first three rows and the last two rows respectively of the Col. 2 of Fig. 7 show the results of style transfer based on real and generated images. In terms of appearance enhancement, our method better maintains the object’s semantic and structure (Row. 1 and 2). In addition, our method can perform certain photographic-level image adjustments, such as background blur and long-exposure effects (Row. 3 and 4), and can also adjust the image’s tones (Row. 5). These impressive results reveal the immense potential of diffusion models in addressing low-level vision tasks.
TABLE I: Quantitative evaluation for stylized images across style transfer and visual enhancement tasks. We report structure distance and human preference scores for our method and other baselines.
Metric Structure Distance $\downarrow$ Human Preference $\uparrow$
Method Ours FreeDiff P2P PNP SDEdit Ours FreeDiff P2P PNP SDEdit
Style Transfer 0.016 0.071 0.053 0.054 0.087 70% 4% 6% 16% 4%
Visual Enhancement 0.011 0.066 0.049 0.041 0.072 78% 6% 8% 8% 0%
TABLE II: CLIP Aesthetic Score Comparisons on Visual Enhancement.
Source Ours FreeDiff P2P PNP SDEdit
CLIP Aes. $\uparrow$ 5.46 6.13 5.37 4.90 5.57 5.25
TABLE III: CLIP Score Comparison on Style Transfer.
Ours FreeDiff P2P PNP SDEdit
CLIP Score $\uparrow$ 0.352 0.346 0.348 0.330 0.329
BG LPIPS $\downarrow$ 0.157 0.160 0.163 0.158 0.177
V-B3 Compare with More Style Transfer Methods
We further qualitatively compare our method with other methods that have different experimental settings, including the diffusion-based fine-tuning methods ControlNet [60], IP-Adapter [57], DiffStyler [25], the text-driven style transfer method based on the CLIP model [37] ClipStyler [29], and the traditional image-driven style transfer methods such as AdaIN [26], AdaAttn [32] and Z* [12]. As shown in Fig. 8, compared to the aforementioned methods, our approach yields more vibrant results in the anime-style transformation task.
Furthermore, for visual enhancement, traditional image-driven style transfer methods AdaIN [26], AdaAttn [32] and Z* fall short as it cannot achieve this effect due to the unavailability of the corresponding reference style image. Simultaneously, our method surpasses text-driven style transfer methods ControlNet [60], IP-Adapter [57], DiffStyler [25], and ClipStyler [29], as it better enhances image’s expressiveness while preserving the structure of the image.
V-C Quantitative Comparison
To better evaluate our method, we use multiple quantitative metrics in ImageNet-R-TI2I [50] dataset for assessment, the results of which are presented in Tab. I. We evaluate the average performance of seven common style transfers, such as “Anime artwork”, “Watercolor painting”, “Play-doh”, “Origami”, “Pixel-art”, and “Cyberpunk” in real-world images for the style transfer task. For visual enhancement in real-world images, we employ five style prompts such as “Cinematic”, “Realistic” and “Best quality”.
As shown in Tab. I, we utilize Structure Distance [49, 36] to measure the structural consistency between the stylized image and the source content image. This measurement is computed by calculating the cosine distance between the two using DINO embeddings [7]. A lower score on Structure Distance means that the structure of the stylized image is more similar to the source content image.
We also present the human preference score in Tab. I. This user study involved a survey of 40 participants, all of whom hold bachelor’s degrees or higher. They were asked to select the best method based on structural consistency and style transfer effects. Each participant was required to evaluate the results of five randomly assigned experimental groups. From the 200 votes collected, it is evident that our method holds significant advantages over other methods.
In addition, we present CLIP Aesthetic Score [42] of visual enhancement on real-world images in Tab. II. A higher CLIP Aesthetic Score indicates that the fusion of style and content is more aesthetically pleasing. We also present the CLIP score in Tab. III to measure the correspondence between style prompts and stylized images. Lastly, to ensure that the background is well preserved after stylization, we report the background LPIPS distance (BG LPIPS) in Tab. III for successfully stylized images. This metric is computed by measuring the LPIPS distance between the background regions of the source and stylized images, identified using a detector. A lower BG LPIPS score indicates better preservation of the source image’s background. The results show the superiority of our method in following textual requirements through stylized image outputs.
@@ V-C Quantitative Comparison
V-D Ablation Study
We conduct ablation experiments on the guidance mask of the AdaBlending module. As shown in Fig. 9 (a)’s anime and neonpunk style trasfer task, noticeable structural drifts are observed when no guidance mask is used. When only $M_{head}$ is used, the overall structure of the image is preserved, but drifts occur in the local details of the image. However, when both spatial and attention-head level masks are used concurrently, AdaBlending delivers consistent structural and semantic style transfer results. These results highlight the crucial role of guidance masks in achieving accurate and consistent style transfer outcomes.
Furthermore, we conducted an ablation study on the hyperparameters $\alpha$ and $n$, varying one while keeping the other constant. As shown in Row. 1 of Fig. 9 (b) for the cinematic transfer task, setting $\alpha=0.7$ achieves effective blurring and tone adjustment while preserving clear semantic content in the image. The Row. 2 of Fig. 9 presents a van Gogh painting transfer task. It can be seen that the stylized image with $n=8$ (without the attention-head mask) exhibits semantic drift compared to the source image. In contrast, when $n=6$, the image has suitable stylistic textural details but lacks color contrast to emphasize the distinct style. However, when $n=7$, the stylized image retains the source image’s structure and achieves effective style transfer.
VI Conclusion
In this work, we aim to study text-driven style transfer. To mitigate the well-known issue of structure inconsistency induced by text-image misalignment in diffusion-based T2I models, we propose a novel and effective tuning-free algorithm, termed Adaptive Style Incorporation (ASI), to incorporate the style information on the fine-grained feature level. It consists of the Siamese Cross-Attention (SiCA) to decouple the single-track cross-attention to a dual-track structure to obtain separate content and style features and the Adaptive Content-Style Blending (AdaBlending) module to couple the content and style information from a structure-consistent manner. We conduct a thorough evaluation of our method across diverse source image domains, both real and generated ones, on a range of common style transfer tasks. Experimental results have demonstrated the effectiveness of our proposed ASI and exhibit significant advantages over SOTA methods.